### v1

Abstract
We observed comet C/2025 A6 with FAST telescope equipped with the ultra-wideband receiver from 23rd October to 8th November 2025 and detected the OH 18-cm lines for the first time. The OH lines underwent a reversal from emission to absorption from 23rd October to 5th November, which is mainly caused by variations in the heliocentric velocity. Through trapezoidal fitting of the OH line profiles, we derive expansion velocities of the water that rise as the heliocentric distance decreases. Based on these results, we estimated the OH production rates of C/2025 A6 for 23rd October, 26th October, 4th November, and 5th November and it presents a significant upward trend.
keywords:
methods: observational– comets: general– comets: individual: C/2025 A61 Introduction
The comets spend most of their lives in the outer, frozen regions of the solar system. Their chemical compositions preserve crucial information about the primordial environment and the processes that may dominated the Solar System’s formation (2017RSPTA.37560261A; 1991ASSL..167.....N). When comets approach the Sun, the solar radiation heats their nuclei, causing subsurface ices to sublimate, and releasing gases and various molecules into the space. As the primary volatile in comets, the production rate and abundance of water provide directly insights into the activity of the nucleus and its physical properties (2004come.book..391B). Under the solar radiation, the water molecules undergo photodissociation: + H2O , producing hydroxyl radicals (OH, 1989A&A...213..459C).
Hydroxyl radicals in cometary coma are primarily observed at ultraviolet and radio wavelengths. At radio frequencies, the OH 18-cm lines arise from -doubling and hyperfine transitions within the ground state of the OH radical, distributed near 1612, 1665, 1667, and 1720 MHz (see Table 1). Compared with UV measurements, OH 18-cm lines can be observed from the ground, and less affected by dust extinction. Their narrow profiles provide direct constraints on the gas expansion velocity and kinematic structure of the coma (1990A&A...238..382B). The OH 18-cm line profiles can be used to measure the expansion velocity of water in the cometary coma (1990A&A...238..382B), providing a basis for further studies of the coma’s gas dynamics. Since the first detection the OH 18-cm lines from Comet Kohoutek in 1973 (1974A&A....34..163B), long-term ground-based observations have established extensive datasets of water and OH emissions for a wide variety of comets (e.g., 1981A&A....99..320D; 2002A&A...393.1053C; 2017AJ....154..249W). These studies have extended observations to comets with different orbital periods and have used variations in molecular production rates to investigate cometary outgassing activity. On the other hand, either the emission or the absorption of the OH 18-cm lines from a comet could be observed, where OH is produced by the reaction between the photon and water molecular, requiring the the ultraviolet light with the wavelength shorter than 186 nm (1989A&A...213..459C).
The Comet C/2025 A6 (Lemmon) is a non-periodic comet. It was discovered by the Mount Lemmon Survey on 3rd January, 2025, and passed closest to Earth on 21st October, 2025, with a geocentric distance of 0.59 AU. It subsequently reached perihelion on 8th November, 2025, at a heliocentric distance of 0.53 AU. Between 3rd January, 2025 and 11th November, 2025, it is in the sky coverage of the Five-hundred-meter Aperture Spherical radio telescope (FAST), which is the largest single-dish radio telescope nowadays in the world. The FAST has been used for comet related studies before (e.g., 2024RAA....24j5008C) with the L-band to search for the molecular spectral lines.
In this work, we present observations of the comet C/2025 A6 with FAST. Section 2 describes the observations. The data processing is mentioned in the Section 3. In Section 4, we discuss the changes of OH line profiles and production rate. The conclusion is presented in Section 5.
| Transitions | Common designation | Frequency (MHz) |
| OH 1612 MHz | 1612.2309 | |
| OH 1665 MHz | 1665.4018 | |
| OH 1667 MHz | 1667.3590 | |
| OH 1720 MHz | 1720.5299 |
2 Observations
The observations of C/2025 A6 were performed with FAST (2011IJMPD..20..989N) on 23rd and 26th October and from 3rd to 8th November in this year. The details can be found in Table 2. It is worth noting that the C/2025 A6 moved very quickly, all observations were conducted in the user-defined mode.
In order to observe the OH 18-cm lines at 1700 MHz, the ultra-wideband (UWB) receiver (2023RAA....23g5016Z) covering 0.5–3.3 GHz was used. This testing receiver is not cooled and with a system temperature between 89 to 130 K. At the spectra frequency (1660 MHz), the beam size is around . A photo of the comet with the size of FAST beam showing that the nuclei can be fully covered (Figure 1). Previous tests have shown excellent agreement between the OH flux densities and velocities measured with the FAST UWB receiver and those obtained with the Arecibo 300 m telescope (2023RAA....23g5016Z).
The data from the UWB receiver were separated into four sub-bands, being 0-1100, 800-1900, 1600-2700 and 2400-3500 MHz, respectively. There are 1,048,576 channels in every sub-band, corresponding to a channel width of 1049 Hz. For OH 18-cm lines, the velocity resolution is approximately 0.19 km s-1. The four polarizations were recorded with an integration time of 1 s. The noise calibration signal was injected with noise diodes for 1 second and stopped for 9 seconds in every 10 seconds during the observation. The noise temperatures of the calibration signal injected to the two polarization channels are 16 and 18 K, respectively.
For fast-moving targets such as the comet, we tracked the comet along its predicted trajectory by updating the telescope pointing every 0.1 s during the observation. In order to find any possible background sources which may affect the comet’s spectra, positions around the comet were also observed in the observation on November 6th. These positions were at least 3′ off the comet nuclei.
| Start Time (UTC) | R.A. (h:m:s) | Dec. (d:m:s) | (AU) | (km s-1) | (AU) | (km s-1) | (K Jy-1) | Total times (s) | Trajectory |
| 2025/10/23 06:30:00 | 14:53:49.44 | +24:36:16.3 | 0.65 | -22.39 | 0.60 | 13.41 | 12.5 | 3000 | Tracking-On |
| 2025/10/26 03:35:00 | 15:31:52.58 | +16:47:47.8 | 0.61 | -19.85 | 0.64 | 29.05 | 12.6 | 10200 | On-Off |
| 2025/11/03 04:44:00 | 16:36:05.31 | -01:25:22.0 | 0.54 | -9.34 | 0.85 | 55.01 | 10.0 | 2400 | Tracking-On |
| 2025/11/04 04:52:00 | 16:40:51.08 | -03:10:07.7 | 0.54 | -7.65 | 0.88 | 56.54 | 10.5 | 4200 | On-Off |
| 2025/11/05 05:01:00 | 16:45:05.35 | -04:48:38.0 | 0.54 | -5.94 | 0.91 | 57.67 | 10.0 | 2760 | On-Off |
| 2025/11/06 05:09:00 | 16:48:50.68 | -06:21:10.8 | 0.53 | -3.01 | 0.97 | 58.89 | 9.5 | 1560 | Multi-point |
| 2025/11/07 05:17:00 | 16:52:09.97 | -07:48:16.1 | 0.53 | -2.37 | 0.98 | 59.10 | 10.0 | 2880 | Tracking-On |
| 2025/11/08 05:27:00 | 16:55:05.90 | -09:10:25.3 | 0.53 | -0.55 | 1.01 | 59.40 | 9.5 | 2220 | Tracking-On |
Notes. The second and third columns list the astrometric right ascension and declination at the beginning of each observation, in the J2000 reference frame. represents the average heliocentric distance during observation, is the average rate of change of the heliocentric distance. is the average distance between the comet and the observatory during the observation, and represents the average rate of change of the comet–observatory distance during the observation. is the gain corresponding to the observation.
3 Data reduction
Due to the limitation of sideband suppression of the UWB, only the bands 500–950, 950–1750, 1750–2550, and 2550–3300 MHz were used (2023RAA....23g5016Z). The OH 18-cm lines are around 1612 to 1720 MHz. Thus, only the spectra in such frequency range were used in our study. For each cycle, the temperature corresponding to the measured power is determined by differencing the spectral line data acquired with open and closed noise diode, followed by calibrating the resulting power values of all spectral lines into temperature. The calibrated spectral data are processed using the astronomical data processing software GILDAS/CLASS111https://www.iram.fr/IRAMFR/GILDAS/. For all observations, a second-order polynomial baseline was fitted and subtracted over a velocity span of 60 km s-1 around the line, excluding the channels containing line emission, absorption features, or identifiable RFI.
The spectra were then smoothed with a two-channel boxcar window, yielding a velocity resolution of about 0.38 km s-1. All spectra were combined with a sigma-weighted averaging.
We performed Gaussian fitting on all spectra with clear detections to obtain the integrated intensity, line-center velocity, FWHM, and peak amplitude of the OH 18-cm lines. We also applied a trapezoidal fitting procedure to the 1667 MHz spectra of the four observations with detections (23rd October, 26th October, 4th November and 5th November) in order to derive the half-width of the trapezoidal base. The derived parameters are summarized in Table 3.
For whom may interested in the conversion from the brightness temperature to the flux with FAST data, the details can be found in appendix A.
| Date (UT) | Lines | RMS (K) | Area (K km s-1) | (K) | S ( Jy km s-1) | (km s-1) | FWHM (km s-1) | (K) | (km s-1) | |
| 2025/10/23 | 1665 | 0.07 | 2.10.1 | 0.45 | 3.3 | 1.70.1 | 13.40.1 | 3.40.2 | 0.59 | 2.40.3 |
| 1667 | 0.07 | 3.90.1 | 3.10.1 | 13.30.1 | 3.10.1 | 1.21 | 2.40.1 | |||
| 2025/10/26 | 1665 | 0.06 | 2.30.1 | 0.51 | 3.5 | 1.80.1 | 28.80.1 | 3.50.2 | 0.61 | 2.90.6 |
| 1667 | 0.06 | 4.80.1 | 3.80.1 | 29.00.1 | 3.40.1 | 1.33 | 2.60.1 | |||
| 2025/11/03 | 1665 | 0.08 | -0.20.1 | -0.27 | 3.1 | -0.20.1 | 54.00.1 | 0.70.3 | -0.23 | - |
| 1667 | 0.09 | -0.40.1 | -0.40.1 | 53.50.1 | 1.10.2 | -0.34 | - | |||
| 2025/11/04 | 1665 | 0.12 | -0.50.1 | -0.37 | 3.1 | -0.50.1 | 55.90.2 | 1.30.3 | -0.38 | - |
| 1667 | 0.09 | -1.40.2 | -1.30.2 | 56.40.2 | 3.20.4 | -0.40 | 3.60.8 | |||
| 2025/11/05 | 1665 | 0.20 | -0.60.2 | -0.33 | 3.1 | -0.70.2 | 57.80.6 | 3.91.1 | -0.16 | - |
| 1667 | 0.09 | -1.30.2 | -1.40.2 | 57.00.3 | 3.90.6 | -0.33 | 3.90.9 |
Notes. Column 2 represents the spectral lines observed in the corresponding frequency range; Column 3 is the 1 noise of the observed spectra with the channel width 0.38 km s-1; Column 4, 8, 9, 10 are derived from Gaussian fitting, representing the integrated intensity, the position of the center of the line obtained through fitting, the full width at half height, and the peak temperature obtained through fitting, respectively; Column 5 is the inversion of ground state of (1988ApJ...331.1058S); Column 7 represents the corresponding flux; Column 11 is the half-length of the base, derived from trapezoidal fitting which represents the sum of the velocities of the parent and daughter molecules.
4 Results and Discussion
During four observing epochs on 23rd October, 26th October, 4th November, and 5th November, either emission or absorption of the 1667 MHz line was detected.
4.1 Change of OH 18-cm Lines from Emission to Absorption
In our observations, the OH 18-cm lines exhibit a clear evolution from emission to absorption. During the first two observing epochs, the OH lines are detected in emission in both the 1665 and 1667 MHz transitions, with corresponding velocities of 13.4 km s-1 and 28.9 km s-1, respectively. In subsequent observations, the lines appear in absorption in both main lines, with a stronger signal at 1667 MHz. The corresponding line-of-sight velocities are 54.0 km s-1, 56.3 km s-1, and 57.0 km s-1 on 3th November, 4th November and 5th November, respectively. The fitted line-center velocities are consistent with the expected radial velocity of the comet relative to the telescope at each observing epoch, confirming the cometary origin of the detected signals. No distinct OH signal is detected during the observations from 6th to 8th November within the expected velocity range.
The observed transition of the OH 18-cm lines from emission to absorption indicates a change in the excitation conditions of the OH ground-state -doublet population. In the cometary coma, these level populations are controlled primarily by resonant solar ultraviolet fluorescence, and the resulting inversion parameter depends on the comet’s heliocentric radial velocity.
The OH band near 3086 Å is pumped by a dense set of intrinsically narrow ro-vibrational transitions. The corresponding Doppler shift can be approximated as
such that at Å, a heliocentric radial velocity of corresponds to a wavelength of Å. At these wavelengths, the solar Fraunhofer spectrum is strongly structured on scales of order 0.1 Å, comparable to the Doppler shifts induced by heliocentric radial velocities of a few km s-1 (see Fig. 2 of 1982ApJ...258..864S).
Consequently, modest heliocentric radial velocity variations during the observing period can alter the OH pumping rate and drive the inversion parameter of the ground-state -doublet across zero. Against the Galactic background continuum, this leads to a transition in the observed OH 18-cm line profiles from emission to absorption (1981A&A....99..320D). Using the inversion curves of 1988ApJ...331.1058S, we derived the inversion parameter at the heliocentric radial velocity of each epoch and compared it with the observed line profiles. The emission detected at early epochs and the absorption observed at later times are consistent with positive () and negative () inversion, respectively. The absence of a clear OH detection in the final observing epochs likely indicates that the inversion parameter was close to zero.
4.2 Water Expansion Velocities from OH Lines
The OH coma is commonly described using a spherically symmetric, steady-state, and isotropic outflow model, in which OH is treated as a secondary product of H2O photodissociation. The velocity of OH molecules is determined by the expansion velocity of the parent gas, , combined with an additional ejection velocity due to photodissociation, (1984A&A...131..111B; 2002A&A...393.1053C). The OH 18-cm lines profiles can be well reproduced by trapezoidal fitting, where the width of the lower base corresponds to 2(+). This trapezoidal profile model has been widely applied in cometary OH line studies (2023A&A...677A.157D; 2025A&A...701A.204L; 2025PSJ.....6..261S). Although OH line profiles always display asymmetries due to anisotropic outgassing toward the Sun or Earth (2025A&A...701A.204L), such effects have only a limited impact on the determination of the lower-base width (1984A&A...131..111B). In this work, we adopt a value of = 0.95 km s-1 to represent the ejection velocity of OH radicals following the photodissociation of H2O, based on the analyses presented in (2002A&A...393.1053C; 2007A&A...467..729T). This value is consistent with the typical OH ejection speeds reported in previous cometary studies (1989A&A...213..459C; 1990A&A...238..382B). The resulting values of derived from the symmetric trapezoidal fits are listed in Table 3.
4.3 The estimate of OH production rate
The OH production rate can be estimated by (1985AJ.....90.1117S; 1990A&A...238..382B):
and
where is total number of OH radicals in the coma, is the integrated intensity of the 1667 MHz line in units of Jy km s-1, is the inversion of the ground state, and is the geocentric distance of the comet in AU. denotes the background temperature in the unit of K, is the OH lifetime, and is correction factor. For , it can be expressed as: = (2017AJ....154..249W; 2025PSJ.....6..261S), and was estimated as s from 1990A&A...238..382B, where the is the heliocentric distance of the comet. Following the method of 2002A&A...393.1053C, we estimated using the 1420 MHz brightness temperature distribution. The 1420 MHz map was obtained from the NASA/LAMBDA222https://lambda.gsfc.nasa.gov/product Stockert–Villa-Elisa all-sky survey (1982A&AS...48..219R; 1986A&AS...63..205R; 2001A&A...368.1123T), and the corresponding background temperature is listed in Table 3. The inversion of ground state at the corresponding coma velocity is obtained from (1988ApJ...331.1058S) and listed in Table 3.
Since the FAST beam is considerably smaller than the characteristic size of the OH coma, a beam-size correction is required to account for flux losses outside the beam. We therefore apply the correction factor to convert the observed number of OH radicals within the beam to the total OH population in the coma. It can be expressed as:
where is the OH column density distribution derived from Haser model. The expansion velocity is derived from the trapezoidal fits, and the photodissociation lifetime of H2O at 1 AU is adopted as s (1992Ap&SS.195....1H), also scaling as . The beam weighting function is assumed to be Gaussian, with a FWHM of 2.5 arcmin.
For small heliocentric distances, the inversion of the OH ground state is quenched by the collision of ions and electrons. Neglecting the contribution of the quenched region will lead to an underestimate of the OH production rate. However, owing to the small FAST beam, which is comparable in size to the quenching region, the iterative quenching correction fails to converge for our observational configuration. Details are provided in Appendix B.
When only the beam correction is applied, the derived values serve as lower limits, with value of , , , and s-1 for 23rd October, 26th October, 4th November, and 5th November, respectively. The production rate basically increases as the heliocentric distance decreases. The water production rate of comet can be approximated as (e.g., 2025PSJ.....6..261S; 2025A&A...701A.204L).
5 Conclusions
We present FAST observations of the OH 18-cm lines in comet C/2025 A6 obtained between 23rd October and 8th November 2025. Our main results and conclusions can be summarized as follows:
1. We confirm that the detected OH 18-cm signals originate from the cometary coma. This conclusion is supported by (i) the agreement between the fitted line-center velocities and radial velocity of the comet relative to the telescope, and (ii) the On–Off observations, in which no emission is detected at the offset positions along the comet’s trajectory. We detected the OH main lines in emission on 23rd and 26th October and in absorption on 3rd to 5th November. The observed emission-to-absorption transition is consistent with the dependence of the OH inversion parameter on heliocentric radial velocity through the Swings effect, which can drive the inversion parameter across zero.
2. We fitted the detected OH line profiles with a trapezoidal model and derived the corresponding parent expansion velocities. The results show that the expansion velocity increases as the heliocentric distance decreases, in agreement with expectations for solar-driven cometary outgassing.
3. Using the measured 1667 MHz line intensities and applying beam-size and quenching corrections, we estimate a plausible range of OH production rates and derive lower limits on . The results show a general increase with decreasing heliocentric distance.
Future observations sampling multiple positions across the coma, or coordinated observations using beams of different sizes, will mitigate systematic uncertainties associated with collisional quenching and provide more robust constraints on .
Acknowledgements.
This work is supported by the National Key R & D Program of China No. 2025SKA0140100, No. 2022YFC2205202, No. 2020SKA0120100, and the National Natural Science Foundation of China (NSFC, Grant Nos. 12373032, 12003047, 11773041, U2031119, 12173052, and 12173053). Both Lei Qian and Zhichen Pan were supported by the Youth Innovation Promotion Association of CAS (id. 2018075, Y2022027, and 2023064) and the CAS “Light of West China” Program. Hongfei Liu has been supported by the National Natural Science Foundation of China (NSFC) under No.12273072. Liyun Zhang has been supported by the Science and Technology Program of Guizhou Province under project No.QKHPTRC-ZDSYS[2023]003 and QKHFQ[2023]003. FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. This work made use of the data from FAST (Five-hundred-meter Aperture Spherical radio Telescope). We thank Prof. Oleg Smirnov and his group at MeerKAT for very helpful discussions. We also acknowledge the support of the University of Chinese Academy of Sciences and the use of its 70 cm telescope.Appendix A From Antenna temperature to Flux for FAST UWB data
The antenna brightness temperature was converted to flux density by using:
Where the G is the antenna gain in the unit of K Jy-1, and is the antenna brightness temperature in the unit of K. The antenna gain can be derived by:
Here, represents the telescope’s efficiency, and is the theoretical gain based on the geometrical illumination area. We adopt the value of the Gain at the 1650 MHz from (2023RAA....23g5016Z), according to the zenith angle (ZA) which they detected derive the corresponding 19.6 K Jy-1, then get the actual gain through the relationship between ZA and (2019SCPMA..6259502J). The result was shown in Table 2.
Appendix B The OH production rate with quenching
When considering the quenching effect, we assume that the inversion factor becomes for , so that OH radicals within this region do not contribute to the observable 18-cm signal (1981A&A....99..320D), and the factor can be written as:
Where the is the column density distribution after subtracted the sphere with radius . The quenching radius can be expressed as = (e.g., 2025A&A...701A.204L). Adopting following 2025A&A...701A.204L results in unstable convergence. We therefore adopt a reduced value of to ensure convergence. With this treatment, the derived values for 23rd October, 26th October, 4th November, and 5th November are , , , and s-1, respectively.